OpenStudy (george.g101):

Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three, solve for f(g(x)) when x = 2. A) 2 B) 5 C) 11 D) 33

OpenStudy (george.g101):

SOS @hartnn @jigglypuff314

OpenStudy (george.g101):

@YanaSidlinskiy @AmTran_Bus

OpenStudy (george.g101):

@sleepyjess @mayankdevnani

OpenStudy (yanasidlinskiy):

\[g(2)= \frac{ 2+4 }{ 3 } = \frac{ 6 }{ 3 } = 2\]

OpenStudy (george.g101):

Ohhh ok, thank you very muchh! :)

OpenStudy (yanasidlinskiy):

\[f(g(2)) = f(2)= (2)^2+ 2\times (2)+3=4+4+3=11\]

OpenStudy (yanasidlinskiy):

Yea no problem:)